OpenStudy (anonymous):

the roots of the polynomial equation 2x^3-8x^2+3x+5=0 are a,b,c. Find the polynomial equation with roots a^2,b^2,c^2. Also, how do you answer these type of questions where they ask you to find different versions of the roots?

mathslover (mathslover):

http://openstudy.com/users/hartnn#/updates/505c45fee4b02e13940f9551 similar to this what have you tried yet @duplicitycheese ?

OpenStudy (anonymous):

here let y=a^2 hence a=sqrt(y) since a is the root of the given equation so a satisfies it hence we r to replace x by sqrt(y) in the above equations and square the terms if(necessary ) to make y free from surds

mathslover (mathslover):

\[\large{a+b+c=0}\]

OpenStudy (anonymous):

Hm, well I've tried fiddling with the product of the roots - squaring it and rearranging it... but it looked too weird

mathslover (mathslover):

@hartnn

hartnn (hartnn):

' hence we r to replace x by sqrt(y)' its correct,show your work

hartnn (hartnn):

but maybe not the best approach here... i can see that becoming complicated

OpenStudy (anonymous):

Oh, I shall try that.

mathslover (mathslover):

good spirit @duplicitycheese

hartnn (hartnn):

hmm...that approach is not difficult.....if u don't get it, ask...

OpenStudy (anonymous):

I got stuck after subbing in sqrt(y). Do I square the whole thing?

hartnn (hartnn):

i will get u to that step..... \(\large 2y^{3/2}-8y+3y^{1/2}+5=0 \\ \large8y-5=y^{1/2}(2y+3)\) Now square both sides.......

OpenStudy (anonymous):

squaring...

OpenStudy (anonymous):

ok, got it but there's still sqrts in here: 4y^3-32y^(5/2)+64y^2=9y+30y^(1/2)+25

hartnn (hartnn):

u did it incorrectly.... (8y-5)^2 = y (2y+3)^2 where does y^(1/2) come from ??

OpenStudy (anonymous):

ah opps, sorry I squared the wrong things.

OpenStudy (anonymous):

Is this the right answer? 4y^3-32^y2+89y-25?

OpenStudy (anonymous):

*52y^2

hartnn (hartnn):

4y^3 -52y^2+89y-25=0

OpenStudy (anonymous):

yay! Thanks!

hartnn (hartnn):

welcome :)